OpenStudy (anonymous):

2 cos x √3 = 0 for trigonometry

OpenStudy (tkhunny):

Notation does not makes sense. Did you mean \(2\cos(x) - \sqrt{3} = 0\)

OpenStudy (anonymous):

If as presented, note that cos x = 0 when x = pi/2, 3 pi/2 etc.

OpenStudy (anonymous):

Sorry it is 2 cos x + √3=0

OpenStudy (anonymous):

So cos x = -sqrt(3)/2 and it looks related to a 30-60-90 triangle except for that minus sign. x = arc cos [-sqrt(3)/2], but perhaps they want you to find the actual angle x.

OpenStudy (anonymous):

There isn't a minus sign that's radical 3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Maybe they are referring to \[A \cos (Bx+c)\]?